# A Two-Timescale Primal-Dual Framework for Reinforcement Learning via Online Dual Variable Guidance

### v2

@@ Appendix C Asynchronous Convergence Rate: Proof Details
$\tilde{\mathcal{P}}_{\rho}^{\,M}(\cdot|s)\ \geq\ |\mathcal{A}|^{M}(\pi_{\min})^{M}\,\mathcal{P}_{d}^{\,M}(\cdot|s)\ \geq\ |\mathcal{A}|^{M}(\pi_{\min})^{M}\,\epsilon_{0}\,\upsilon(\cdot)\ =:\ \epsilon\,\upsilon(\cdot),$
uniformly in $\rho$ and $s$. Thus each $\tilde{\mathcal{P}}_{\rho}^{\,M}$ satisfies a Doeblin minorization with the same $(\epsilon,\upsilon)$, implying the Dobrushin coefficient $\delta(\tilde{\mathcal{P}}_{\rho}^{\,M})\leq 1-\epsilon$ for all $\rho$. For $k=mM+r$ with $0\leq r<M$, that is, after $m$ blocks of length $M$, and any probability measures $\mu,\mu^{\prime}$ on $\mathcal{X}$, by the sub-multiplication property of the Dobrushin coefficient and the non-expansivity of stochastic kernels in total variation,
$\|\mu\,\mathcal{P}_{\rho}^{k}-\mu^{\prime}\,\mathcal{P}_{\rho}^{k}\|_{\mathrm{TV}}\ \leq\ (1-\epsilon)^{m}\,\|\mu-\mu^{\prime}\|_{\mathrm{TV}}\ \leq\,(1-\epsilon)^{\frac{k-(M-1)}{M}}\ =\ \varrho^{-(M-1)}\varrho^{k},$
for $\varrho:=(1-|\mathcal{A}|^{M}(\pi_{\min})^{M}\epsilon_{0})^{1/M}$, where in the second inequality we used $\|\mu-\mu^{\prime}\|_{\mathrm{TV}}\leq 1$. Setting $\mu^{\prime}=\mu_{\rho}$, the stationary distribution of $\mathcal{P}_{\rho}$ that exists as a consequence of the Doeblin minorization, and with $C_{\mathcal{X}}:=\varrho^{-(M-1)}$, this gives the uniform geometric ergodicity constants independent of $\rho\in H$. $\blacksquare$
@@
Next, we provide the proofs of the convergence rate-related properties of the iterates generated by Algorithm 2 with primal variable projection to $\mathcal{V}_{r}$, summarized in Proposition 4.2. The visitation floor and replay-buffer bias decay results are of particular interest since they are the main differences of our asynchronous approach compared to Zeng et al. [2024].
@@
We start with the visitation floor property, stated in Proposition 4.2(i).
Lemma C.1.
@@
(ii) To show the high-probability linear growth bound for all state-action pairs, we first establish lower bounds for the time-varying expected visits, then apply a concentration bound for inhomogeneous Markov chains by Paulin [2015] to establish concentration for any time step $k$, and lastly apply a union bound to achieve the high-probability minimal growth bound for all steps $k\geq K(\delta)$.
We begin by deriving a block contraction of the time-inhomogeneous kernel sequence with Dobrushin’s ergodic coefficient [Dobrushin, 1956], [Brémaud, 2020, Definition 4.3.11]. For a Markov chain with transition kernel $Q$ the ergodic coefficient is defined as
By [Brémaud, 2020, Corollary 4.3.17] the coefficient satisfies the contraction inequality $\|\mu Q-\nu Q\|_{\mathrm{TV}}\leq\delta(Q)\|\mu-\nu\|_{\mathrm{TV}}$ for any two probability measures $\mu,\ \nu$ on $\mathcal{X}$.
@@
Under Assumption 3, the proof of Lemma 4.1 shows that we can choose constants $m_{\star}\in\mathbb{N}$ and $\kappa\in(0,1)$ such that
@@
The stationary distributions are Lipschitz continuous in the dual variable (see Lemma C.7), so there exists $L_{\mu}>0$ such that $\|\mu_{\rho_{t}}-\mu_{\rho_{t+1}}\|_{\mathrm{TV}}\ \leq\ L_{\mu}\|\rho_{t+1}-\rho_{t}\|_{2}.$ Therefore,
@@
and we can apply the one-sided McDiarmid inequality for time-inhomogeneous Markov chains [Paulin, 2015, Corollary 2.10] to $\varphi$ and obtain
@@
see [Paulin, 2015, Definition 1.4], is the inhomogeneous mixing time for $\{\mathcal{P}_{\rho_{k}}\}$. It holds that $\tau_{\min}<\infty$ with a simple bound in our setting deferred to Lemma C.6. From the mean lower bound in (40), there exists $K_{0}\big(\tfrac{p_{\star}}{2}\big)$ such that, for all $k\geq K_{0}\big(\tfrac{p_{\star}}{2}\big)$,
@@
Next, we prove the replay-buffer bias decay stated in Proposition 4.2(ii). In the following, we use the shorthand notation $\bar{\mathcal{E}}_{k}(\rho)$ for the stationary law expectation of the bias vector given a fixed $\rho\in H$, that is,
@@
Lemma C.2 (Replay-buffer bias under high-probability visitation).
Fix $\delta\in(0,1)$, let $\mathcal{G}_{\delta}$ denote the high-probability event of Lemma C.1, and let $K(\delta)$ be the corresponding burn-in time. Define
@@
Proof For each fixed $(s,a)\in\mathcal{X}$ and corresponding buffer size $|\mathcal{D}_{k}(s,a)|=\nu_{k}(s,a)$, the $L^{1}$-bound for a categorical law on $|\mathcal{S}|$ outcomes [Weissman et al., 2003, Theorem 2.1] gives for all $\epsilon>0$
@@
On the high-probability visitation event $\mathcal{G}_{\delta}$ of Lemma C.1, we have $\underline{\nu}_{k}\geq\tfrac{p_{\star}}{2}\,k$ for all $k\geq K(\delta)$, hence
@@
We next address the regularity properties stated in Proposition 4.2(iii). The $\eta_{V}$-strong convexity of the mapping $V\mapsto L(V,\rho)$ has been established already in equations (32) and the proof of Proposition 3.2. The explicit Lipschitz constant of the best response mapping $\lambda$ is derived in the following.
Lemma C.3.
@@
Proof.
The displayed linear form (47) of the optimizer follows from the first-order optimality condition derived in the proof of Proposition 3.2. For the Lipschitz constant, we define the matrix $M\in\mathbb{R}^{|\mathcal{S}||\mathcal{A}|\times|\mathcal{S}|}$ with rows
@@
Then $\|\lambda(\rho)-\lambda(\hat{\rho})\|_{2}=\frac{1}{\eta_{V}}\|M^{\top}(\rho-\hat{\rho})\|_{2}\leq\frac{\|M^{\top}\|_{2}}{\eta_{V}}\|\rho-\hat{\rho}\|_{2}$. Each row of $M$ satisfies $\|\text{row}_{sa}(M)\|_{2}^{2}\leq\|e^{\top}_{s}\|_{2}^{2}+\gamma^{2}\|\mathcal{P}(\cdot|s,a)\|_{2}^{2}\leq(1+\gamma^{2})$. Therefore, the claim follows by noting $\|M^{\top}\|_{2}^{2}=\|M\|_{2}^{2}\leq\|M\|_{F}^{2}=\sum_{(s,a)\in\mathcal{S}\times\mathcal{A}}\|\text{row}_{sa}(M)\|_{2}^{2}\leq|\mathcal{S}||\mathcal{A}|(1+\gamma^{2})$. ∎
@@
since $\Delta[V](s,a)=(-MV)(s,a)+r(s,a)$ and hence, $L(V,\rho)=\frac{\eta_{V}}{2}\|V\|^{2}_{2}-\langle\rho,MV\rangle+\langle\rho,r\rangle+\eta_{\rho}g(\rho)$. Then $\nabla f(\rho)=-\frac{MM^{\top}}{\eta_{V}}\rho+r+\eta_{\rho}\nabla g(\rho)$ and we obtain the following smoothness for the reduced objective
$\|\nabla f(\rho)-\nabla f(\hat{\rho})\|_{2}\ \leq\Big(\eta_{\rho}\frac{2}{C^{L}}+\frac{\|M\|^{2}_{2}}{\eta_{V}}\Big)\|\rho-\hat{\rho}\|_{2}\leq\Big(\eta_{\rho}\frac{2}{C^{L}}+\frac{|\mathcal{S}||\mathcal{A}|(1+\gamma^{2})}{\eta_{V}}\Big)\|\rho-\hat{\rho}\|_{2},\ \forall\,\rho,\hat{\rho}\in H.$
where $\tfrac{2}{C^{L}}$ is the Lipschitz constant of $\nabla g(\rho)$ on H as derived in the following Lemma, and where we have used the upper bound for $\|M\|_{2}^{2}$ from the proof of Lemma C.3.
Lemma C.4 (Lipschitz gradient of $g$).
@@
Proof.
The coordinate-wise gradient is
@@
Lemma C.5 (Strong concavity of the reduced objective).
The reduced objective $f(\rho)=\min_{V}L(V,\rho)$ is $\mu$-strongly concave on $H$ with any modulus $0<\mu\leq\mu_{\mathrm{opt}}$, where
@@
Proof.
As shown previously, it holds that
@@
For any $h\in\mathbb{R}^{|\mathcal{S}||\mathcal{A}|}$, decompose $h=u+v$ per state, where for each $s\in\mathcal{S}$, $\sum_{a\in\mathcal{A}}u(s,a)=0$ and $v(s,a)=q(s)\,\frac{\rho(s,a)}{\tilde{\rho}(s)}$ for some $q(s)\in\mathbb{R}$. With this decomposition, $u$ redistributes mass across actions while keeping the state marginal fixed, and $v$ changes only the state marginals. In particular, $v^{\top}\nabla^{2}g(\rho)v=0$. For any $u$ with $\sum_{a\in\mathcal{A}}u(s,a)=0$, it holds that for all $s\in\mathcal{S}$ and corresponding block matrices $H_{s}(\rho)$ in the Hessian of $g$, specified in the proof of Lemma C.4,
@@
It holds that $\|M^{\top}u\|_{2}\leq\|M\|_{2}\|u\|_{2}$, and together with the operator bound $\|M\|^{2}_{2}\leq|\mathcal{S}||\mathcal{A}|(1+\gamma^{2})$ derived in the proof of Lemma C.3, we have
@@
Since $\sum_{a\in\mathcal{A}}v(s,a)=q(s)$ and $\sum_{a\in\mathcal{A}}v(s,a)\mathcal{P}(\cdot|s,a)=q(s)\sum_{a\in\mathcal{A}}\pi_{\rho}(a|s)\mathcal{P}(\cdot|s,a)=q(s)\bar{\mathcal{P}}_{\rho}(\cdot|s)$,
@@
Since $A>0$ and $C>0$ by construction, the discriminant $(A+B+C)^{2}-4AC$ is strictly smaller than $(A+B+C)^{2}$, so the bracket is positive and hence $\mu_{\mathrm{opt}}>0$. Thus $\nabla^{2}f(\rho)\succeq\mu I$ on $H$ for every $0<\mu\leq\mu_{\mathrm{opt}}$, which proves strong concavity. ∎
@@
Then for all attainable $(\rho,V,X,\Xi)$ it holds that $\|\hat{g}(\rho,V,X,\Xi)\|_{2}\leq D$, $\|\nabla_{V}L(V,\rho)\|_{2}\leq D$, and $\|\nabla_{\rho}L(V,\rho)\|_{2}\leq B$. This completes the proof of Proposition 4.2.
C.2 Proofs of the Propositions 4.3 and 4.4
@@
and we omit the dependence of the bias term under stationary-law (see (44)) on $\rho_{k-1}$ by setting
@@
$\penalty 10000\ \penalty 10000\ \penalty 10000\ +\bar{\alpha}_{k+1}^{2}\,\mathbb{E}\big[\|\widehat{G}_{k+1}\|_{2}^{2}|\mathcal{F}_{k}^{-}\big].$
@@
(III.) The technical proof of the following bound on $2\bar{\alpha}_{k+1}\langle V_{k}-\lambda(\rho_{k}),\Delta^{X}_{k+1}\rangle$ applies a comparison of a frozen-chain to the mixing of the inhomogeneous chain and is provided in Lemma C.8. We obtain the following
@@
where the last inequality uses the bound on the bias under stationary law derived in Lemma C.2.
@@
$\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ +C_{1}\tau_{k}^{2}\alpha_{\mathrm{env},k-\tau_{k}}\alpha_{\mathrm{env},k}(\mathbb{E}[\|V_{k-\tau_{k}}-\lambda(\rho_{k-\tau_{k}})\|_{2}^{2}|\mathcal{G}_{\delta}]+1)$
$\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ +\tfrac{C_{\mathrm{buf}}(\delta)^{2}|\mathcal{S}|}{\tilde{\eta}_{V}}\tfrac{\alpha_{\mathrm{env},k}}{k}+D^{2}\alpha_{\mathrm{env},k}^{2}$
@@
With the Lipschitz continuity of the best response (see Lemma C.3) and the bound on the dual update, we have
@@
$\penalty 10000\ +(1+\tfrac{\tilde{\eta}_{V}}{2}\alpha_{k})\tfrac{C_{\mathrm{buf}}(\delta)^{2}|\mathcal{S}|}{\tilde{\eta}_{V}}\tfrac{\alpha_{\mathrm{env},k}}{k}+(1+\tfrac{\tilde{\eta}_{V}}{2}\alpha_{k})D^{2}\alpha_{\mathrm{env},k}^{2}$
$\penalty 10000\ +L^{2}B^{2}\beta_{\mathrm{env},k}^{2}+\tfrac{2L^{2}B^{2}\beta_{\mathrm{env},k}^{2}}{\tilde{\eta}_{V}\,\alpha_{k}}$
@@
$\penalty 10000\ +\tfrac{9}{2}C_{1}\tau_{k}^{2}\alpha_{\mathrm{env},k-\tau_{k}}\alpha_{\mathrm{env},k}(D^{2}\tau_{k}^{2}\alpha_{\mathrm{env},k-\tau_{k}}^{2}+L^{2}B^{2}\tau_{k}^{2}\beta_{\mathrm{env},k-\tau_{k}}^{2}+\tfrac{1}{3})$
$\penalty 10000\ +\tfrac{3}{2}\tfrac{C_{\mathrm{buf}}(\delta)^{2}|\mathcal{S}|}{\tilde{\eta}_{V}}\tfrac{\alpha_{\mathrm{env},k}}{k}+\tfrac{3}{2}D^{2}\alpha_{\mathrm{env},k}^{2}+L^{2}B^{2}\beta_{\mathrm{env},k}^{2}+\tfrac{2L^{2}B^{2}\beta_{\mathrm{env},k}^{2}}{\tilde{\eta}_{V}\,\alpha_{k}}$
@@
$\penalty 10000\ +\tfrac{3}{2}D^{2}\alpha_{\mathrm{env},k}^{2}+L^{2}B^{2}\beta_{\mathrm{env},k}^{2}+\tfrac{2L^{2}B^{2}}{\tilde{\eta}_{V}}\tfrac{\beta_{\mathrm{env},k}^{2}}{\alpha_{k}},$

### v3

@@ C.1 Proofs Related to Assumption 13 and Proposition 15
$\prod_{t=i}^{i+M-1}\tilde{\mathcal{P}}_{\rho_{t}}(\cdot|s)\geq\pi_{\min}^{M}\mathcal{P}_{d}^{\,M}(\cdot|s)\geq\epsilon\,\upsilon(\cdot),\qquad\forall s\in\mathcal{S}.$
Hence,
$\delta\!\left(\prod_{t=i}^{i+M-1}\tilde{\mathcal{P}}_{\rho_{t}}\right)\leq 1-\epsilon.$
Now using again the factorization of the state-action kernels, we have
$\prod_{t=i}^{i+M}\mathcal{P}_{\rho_{t}}=\mathcal{T}\left(\prod_{t=i}^{i+M-1}\tilde{\mathcal{P}}_{\rho_{t}}\right)\mathcal{B}_{\rho_{i+M}}.$
Therefore, for any probability measures $\mu,\mu^{\prime}$ on $\mathcal{X}$,
$\left\|\mu\prod_{t=i}^{i+M}\mathcal{P}_{\rho_{t}}-\mu^{\prime}\prod_{t=i}^{i+M}\mathcal{P}_{\rho_{t}}\right\|_{\mathrm{TV}}$ $\leq\left\|(\mu\mathcal{T})\prod_{t=i}^{i+M-1}\tilde{\mathcal{P}}_{\rho_{t}}-(\mu^{\prime}\mathcal{T})\prod_{t=i}^{i+M-1}\tilde{\mathcal{P}}_{\rho_{t}}\right\|_{\mathrm{TV}}$
$\leq(1-\epsilon)\|\mu\mathcal{T}-\mu^{\prime}\mathcal{T}\|_{\mathrm{TV}}$
$\leq(1-\epsilon)\|\mu-\mu^{\prime}\|_{\mathrm{TV}}.$
Taking the supremum over $\mu,\mu^{\prime}$ yields
$\delta\!\left(\prod_{t=i}^{i+M}\mathcal{P}_{\rho_{t}}\right)\leq 1-\epsilon.$
Since $i$ was arbitrary, Assumption 13(ii) holds with
$m_{\star}=M+1,\qquad\kappa=1-\epsilon.$
This completes the proof. $\blacksquare$
@@
Next, we provide the proofs of the convergence rate-related properties of the iterates generated by Algorithm 2 with primal variable projection to $\mathcal{V}_{r}$, summarized in Proposition 15. The visitation floor and replay-buffer bias decay results are of particular interest since they are the main differences of our asynchronous approach compared to Zeng et al. (2024).
@@
We start with the visitation floor property, stated in Proposition 15(i).
Lemma 25
@@
(ii) To show the high-probability linear growth bound for all state-action pairs, we first establish lower bounds for the time-varying expected visits, then apply a concentration bound for inhomogeneous Markov chains by Paulin (2015) to establish concentration for any time step $k$, and lastly apply a union bound to achieve the high-probability minimal growth bound for all steps $k\geq K(\delta)$.
We begin by deriving a block contraction of the time-inhomogeneous kernel sequence with Dobrushin’s ergodic coefficient.
By (Brémaud, 2020, Corollary 4.3.17) the coefficient satisfies the contraction inequality $\|\mu Q-\nu Q\|_{\mathrm{TV}}\leq\delta(Q)\|\mu-\nu\|_{\mathrm{TV}}$ for any two probability measures $\mu,\ \nu$ on $\mathcal{X}$.
@@
Under Assumption 13, we can choose constants $m_{\star}\in\mathbb{N}$ and $\kappa\in(0,1)$ such that
@@
The stationary distributions are Lipschitz continuous in the dual variable (see Lemma 32), so there exists $L_{\mu}>0$ such that $\|\mu_{\rho_{t}}-\mu_{\rho_{t+1}}\|_{\mathrm{TV}}\ \leq\ L_{\mu}\|\rho_{t+1}-\rho_{t}\|_{2}.$ Therefore,
@@
and we can apply the one-sided McDiarmid inequality for time-inhomogeneous Markov chains (Paulin, 2015, Corollary 2.10) to $\varphi$ and obtain
@@
see (Paulin, 2015, Definition 1.4), is the inhomogeneous mixing time for $\{\mathcal{P}_{\rho_{k}}\}$. It holds that $\tau_{\min}<\infty$ with a simple bound in our setting deferred to Lemma 31. From the mean lower bound in (40), there exists $K_{0}\big(\tfrac{p_{\star}}{2}\big)$ such that, for all $k\geq K_{0}\big(\tfrac{p_{\star}}{2}\big)$,
@@
Next, we prove the replay-buffer bias decay stated in Proposition 15(ii). In the following, we use the shorthand notation $\bar{\mathcal{E}}_{k}(\rho)$ for the stationary law expectation of the bias vector given a fixed $\rho\in H$, that is,
@@
Lemma 26 (Replay-buffer bias under high-probability visitation)
Fix $\delta\in(0,1)$, let $\mathcal{G}_{\delta}$ denote the high-probability event of Lemma 25, and let $K(\delta)$ be the corresponding burn-in time. Define
@@
Proof  For each fixed $(s,a)\in\mathcal{X}$ and corresponding buffer size $|\mathcal{D}_{k}(s,a)|=\nu_{k}(s,a)$, the $L^{1}$-bound for a categorical law on $|\mathcal{S}|$ outcomes (Weissman et al., 2003, Theorem 2.1) gives for all $\epsilon>0$
@@
On the high-probability visitation event $\mathcal{G}_{\delta}$ of Lemma 25, we have $\underline{\nu}_{k}\geq\tfrac{p_{\star}}{2}\,k$ for all $k\geq K(\delta)$, hence
@@
We next address the regularity properties stated in Proposition 15(iii). The $\eta_{V}$-strong convexity of the mapping $V\mapsto L(V,\rho)$ has been established already in equations (32) and the proof of Proposition 8. The explicit Lipschitz constant of the best response mapping $\lambda$ is derived in the following.
Lemma 27
@@
Proof  The displayed linear form (47) of the optimizer follows from the first-order optimality condition derived in the proof of Proposition 8. For the Lipschitz constant, we define the matrix $M\in\mathbb{R}^{|\mathcal{S}||\mathcal{A}|\times|\mathcal{S}|}$ with rows
@@
Then $\|\lambda(\rho)-\lambda(\hat{\rho})\|_{2}=\frac{1}{\eta_{V}}\|M^{\top}(\rho-\hat{\rho})\|_{2}\leq\frac{\|M^{\top}\|_{2}}{\eta_{V}}\|\rho-\hat{\rho}\|_{2}$. Each row of $M$ satisfies $\|\text{row}_{sa}(M)\|_{2}^{2}\leq\|e^{\top}_{s}\|_{2}^{2}+\gamma^{2}\|\mathcal{P}(\cdot|s,a)\|_{2}^{2}\leq(1+\gamma^{2})$. Therefore, the claim follows by noting $\|M^{\top}\|_{2}^{2}=\|M\|_{2}^{2}\leq\|M\|_{F}^{2}=\sum_{(s,a)\in\mathcal{S}\times\mathcal{A}}\|\text{row}_{sa}(M)\|_{2}^{2}\leq|\mathcal{S}||\mathcal{A}|(1+\gamma^{2})$. $\blacksquare$
@@
[…]
$\|\nabla f(\rho)-\nabla f(\hat{\rho})\|_{2}\ \leq\Big(\eta_{\rho}\frac{2}{C^{L}}+\frac{\|M\|^{2}_{2}}{\eta_{V}}\Big)\|\rho-\hat{\rho}\|_{2}\leq\Big(\eta_{\rho}\frac{2}{C^{L}}+\frac{|\mathcal{S}||\mathcal{A}|(1+\gamma^{2})}{\eta_{V}}\Big)\|\rho-\hat{\rho}\|_{2},\$
where $\tfrac{2}{C^{L}}$ is the Lipschitz constant of $\nabla g(\rho)$ on H as derived in the following Lemma, and where we have used the upper bound for $\|M\|_{2}^{2}$ from the proof of Lemma 27.
Lemma 28 (Lipschitz gradient of $g$)
@@
Proof  The coordinate-wise gradient is
@@
Lemma 29 (Strong concavity of the reduced objective)
The reduced objective
$f(\rho)=\min_{V}L(V,\rho)$ is $\mu$-strongly concave on $H$ with any modulus $0<\mu\leq\mu_{\mathrm{opt}}$, where
@@
Proof  As shown previously, it holds that
@@
For any $h\in\mathbb{R}^{|\mathcal{S}||\mathcal{A}|}$, decompose $h=u+v$ per state, where for each $s\in\mathcal{S}$, $\sum_{a\in\mathcal{A}}u(s,a)=0$ and $v(s,a)=q(s)\,\frac{\rho(s,a)}{\tilde{\rho}(s)}$ for some $q(s)\in\mathbb{R}$. With this decomposition, $u$ redistributes mass across actions while keeping the state marginal fixed, and $v$ changes only the state marginals. In particular, $v^{\top}\nabla^{2}g(\rho)v=0$. For any $u$ with $\sum_{a\in\mathcal{A}}u(s,a)=0$, it holds that for all $s\in\mathcal{S}$ and corresponding block matrices $H_{s}(\rho)$ in the Hessian of $g$, specified in the proof of Lemma 28,
@@
It holds that $\|M^{\top}u\|_{2}\leq\|M\|_{2}\|u\|_{2}$, and together with the operator bound $\|M\|^{2}_{2}\leq|\mathcal{S}||\mathcal{A}|(1+\gamma^{2})$ derived in the proof of Lemma 27, we have
@@
Since $\sum_{a\in\mathcal{A}}v(s,a)=q(s)$ and
$\sum_{a\in\mathcal{A}}v(s,a)\mathcal{P}(\cdot|s,a)=q(s)\sum_{a\in\mathcal{A}}\pi_{\rho}(a|s)\mathcal{P}(\cdot|s,a)=q(s)\bar{\mathcal{P}}_{\rho}(\cdot|s),$
we have
@@
Since $A>0$ and $C>0$ by construction, the discriminant $(A+B+C)^{2}-4AC$ is strictly smaller than $(A+B+C)^{2}$, so the bracket is positive and hence $\mu_{\mathrm{opt}}>0$. Thus $\nabla^{2}f(\rho)\succeq\mu I$ on $H$ for every $0<\mu\leq\mu_{\mathrm{opt}}$, which proves strong concavity. $\blacksquare$
@@
Then for all attainable $(\rho,V,X,\Xi)$ it holds that $\|\hat{g}(\rho,V,X,\Xi)\|_{2}\leq D$, $\|\nabla_{V}L(V,\rho)\|_{2}\leq D$, and $\|\nabla_{\rho}L(V,\rho)\|_{2}\leq B$. This completes the proof of Proposition 15.
C.2 Proofs of the Propositions 18 and 19
@@
and we omit the dependence of the bias term under stationary-law (see 44) on $\rho_{k-1}$ by setting
@@
$~~~+\bar{\alpha}_{k+1}^{2}\,\mathbb{E}\big[\|\widehat{G}_{k+1}\|_{2}^{2}|\mathcal{F}_{k}^{-}\big].$
@@
(III.) The technical proof of the following bound on $2\bar{\alpha}_{k+1}\langle V_{k}-\lambda(\rho_{k}),\Delta^{X}_{k+1}\rangle$ applies a comparison of a frozen-chain to the mixing of the inhomogeneous chain and is provided in Lemma 33. We obtain the following
@@
where the last inequality uses the bound on the bias under stationary law derived in Lemma 26.
@@
$~~~~~+C_{1}\tau_{k}^{2}\alpha_{\mathrm{env},k-\tau_{k}}\alpha_{\mathrm{env},k}(\mathbb{E}[\|V_{k-\tau_{k}}-\lambda(\rho_{k-\tau_{k}})\|_{2}^{2}|\mathcal{G}_{\delta}]+1)$
$~~~~~+\tfrac{C_{\mathrm{buf}}(\delta)^{2}|\mathcal{S}|}{\tilde{\eta}_{V}}\tfrac{\alpha_{\mathrm{env},k}}{k}+D^{2}\alpha_{\mathrm{env},k}^{2}$
@@
With the Lipschitz continuity of the best response (see Lemma 27) and the bound on the dual update, we have
@@
$~+(1+\tfrac{\tilde{\eta}_{V}}{2}\alpha_{k})\tfrac{C_{\mathrm{buf}}(\delta)^{2}|\mathcal{S}|}{\tilde{\eta}_{V}}\tfrac{\alpha_{\mathrm{env},k}}{k}+(1+\tfrac{\tilde{\eta}_{V}}{2}\alpha_{k})D^{2}\alpha_{\mathrm{env},k}^{2}$
$~+L^{2}B^{2}\beta_{\mathrm{env},k}^{2}+\tfrac{2L^{2}B^{2}\beta_{\mathrm{env},k}^{2}}{\tilde{\eta}_{V}\,\alpha_{k}}$
@@
$~+\tfrac{9}{2}C_{1}\tau_{k}^{2}\alpha_{\mathrm{env},k-\tau_{k}}\alpha_{\mathrm{env},k}(D^{2}\tau_{k}^{2}\alpha_{\mathrm{env},k-\tau_{k}}^{2}+L^{2}B^{2}\tau_{k}^{2}\beta_{\mathrm{env},k-\tau_{k}}^{2}+\tfrac{1}{3})$
$~+\tfrac{3}{2}\tfrac{C_{\mathrm{buf}}(\delta)^{2}|\mathcal{S}|}{\tilde{\eta}_{V}}\tfrac{\alpha_{\mathrm{env},k}}{k}+\tfrac{3}{2}D^{2}\alpha_{\mathrm{env},k}^{2}+L^{2}B^{2}\beta_{\mathrm{env},k}^{2}+\tfrac{2L^{2}B^{2}\beta_{\mathrm{env},k}^{2}}{\tilde{\eta}_{V}\,\alpha_{k}}$
@@
$~+\tfrac{3}{2}D^{2}\alpha_{\mathrm{env},k}^{2}+L^{2}B^{2}\beta_{\mathrm{env},k}^{2}+\tfrac{2L^{2}B^{2}}{\tilde{\eta}_{V}}\tfrac{\beta_{\mathrm{env},k}^{2}}{\alpha_{k}},$